The time period of simple pendulum is 2s.what Will be it's length on earth



1
Expert's answer
2022-01-05T11:47:33-0500

The length can be found from the period:


"T=2\pi\sqrt{\dfrac lg}→l=\dfrac{gT^2}{4\pi^2}=0.99\text{ m}."